OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)? -3/2 -2/3 3/2 2/3

OpenStudy (harsimran_hs4):

given two points x1, y1 and x2, y2 the slope of the line joining them is m = (y1-y2) /(x1-x2)

OpenStudy (anonymous):

its -1.5 but i dont kno how to turn that into a fraction

OpenStudy (harsimran_hs4):

-1.5 = -15/10 = -3/2